OpenStudy (anonymous):

19 ÷ 8 = 2 R3 make it a fraction

OpenStudy (accessdenied):

When we divide two numbers and end up with a remainder, we make it a fraction by taking the remainder over the divisor, 8. \[ 2\frac{3}{8} \] We can make this an improper fraction by multiplying the 2 by 8 and adding it to the numerator. \[ 2\frac{3}{8} = \frac{19}{8} \] This actually is what we had originally, only we now have a fraction for it

OpenStudy (anonymous):

17 ÷ 3 = 5 R2 what about this one

OpenStudy (accessdenied):

Same as before, we have the whole number 5 and take the remainder over divisor... \[ 5\frac{2}{3} \] This is the mixed number, then we multiply denominator by 5 and add to 2 for improper fraction. \[ 5\frac{2}{3} = \frac{3*5 + 2}{3} \]

OpenStudy (anonymous):

thanks SOOOOOOOOOO MUCH

OpenStudy (accessdenied):

No problem! :)